I have read this question over and over and I just don't understand it, help please!
The sum of the page numbers on facing pages of a book is 157; what are the pages numbers?

Answer by checkley77(12844)

X+(X+1)=157
2X=157-1
2X=156
X=156/2
X=78 IS THE LEFT HAND PAGE NUMBER.
78+1=79 IS THE NUMBER OF THE RIGHT HAND PAGE.